# On automatic boundedness of operators between ordered and topological vector spaces

### 11 Mar 2026 v2

@@ Abstract
We study topological boundedness of order-to-topology bounded and order-to-topology continuous operators from ordered vector spaces to topological vector spaces. Levi and Lebesgue operators are investigated.
@@ 11 Sobolev Institute of Mathematics, Novosibirsk, Russia
22 Southern Mathematical Institute VSC RAS, Vladikavkaz, Russia
Keywords: ordered vector space, topological vector space, ordered Banach space, ordered Fréchet space, order-to-topology bounded operator
@@ 1 Introduction and preliminaries
Automatic boundedness of operators (enjoying some additional properties) plays significant role in analysis. For example, a nice theory exists on automatic continuity of algebra homomorphisms (see, e.g., [8, 9]). Another classical case is the automatic continuity of positive operators from Banach lattices to normed lattices [3, 4, 7]. Apparently, the importance of boundedness of operators from a Banach space to a normed space lies in the Uniform Boundedness Principle for families of such operators. The present paper is devoted to the question how far one can go with several other kinds of operators from a Banach space to a topological vector space. This question has attracted an attention in several recent papers (see, for example, [5, 14, 15, 13, 16, 20, 18, 19, 22, 23, 25]).
Throughout, vector spaces are real, operators are linear, ${\cal L}(X,Y)$ stands for the space of operators between vector spaces $X$ and $Y$. We abbreviate a topological vector space (normed space, locally convex space, ordered vector space, ordered normed space, ordered Banach space) by TVS (NS, LCS, OVS, ONS, OBS). Symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net such that $\inf\limits_{\alpha}x_{\alpha}=0$, and $a,b]=\{x\in X:a\leq x\leq b\}$ for the order interval in an OVS $X$, where $a,b\in X$. Throughout, the symbol ${\cal L}(X,Y)$ stands for the space of operators between vector spaces $X$ and $Y$, and ${\rm L}(X,Y)$ for the space of bounded operators between NSs $X$ and $Y$, $B_{X}(a,r)=\{x\in X:d(x,a)\leq r\}$ for the closed ball centered at $a\in X$ of radius $r\geq 0$ in a metric space $(X,d)$, and $B_{X}$ for $B_{X}(0,1)$ in NS $X$. It is proved in [15, Theorem 2.1] that order-to-norm bounded families of operators from a Banach lattice to an NS are equi-continuous. In [16, Theorem 2.1], it was extended to the OBS setting.
Let $X$ be a Banach space and $(Y,\tau)$ be a TVS. In absence of an additional structure in $X$, one can not say much about conditions providing boundedness of an operator $T:X\to Y$. Using of finite sets gives no proper subclass, because $T$ always carries finite sets into bounded ones. Whereas, the compactness returns us to the class of bounded operators, namely $T$ takes compact subsets into bounded sets iff $T$ is bounded. To see this, assume that $T$ is not bounded. Then, there exist an absorbing neighborhood $U\in\tau(0)$ and a sequence $(x_{n})$ in $B_{X}$ with $Tx_{n}\notin n^{2}U$ for all $n$. So, the image $\big\{T(\frac{x_{n}}{n})\big\}_{n=1}^{\infty}\cup\{0\}$ of the compact set $\big\{\frac{x_{n}}{n}\big\}_{n=1}^{\infty}\cup\{0\}$ is not bounded.
The case of a Hilbert space in the domain was investigated in papers [18, 19], where it is proved in [19, Theorem 2.1] (see also [18, Lemma 1.2]) that
@@
Let us drop requirements of additional structures (like lattice operations, order, or inner products) on a space in the domain and consider:
@@
The ‘‘conditional part’’ of Question 1.1 is open even when $X$ is reflexive, whereas the ‘‘unconditional part’’ has a negative answer due to Gowers and Maurey, who constructed a reflexive Banach space with no unconditional basic sequence [21]. For such a Banach space $X$, every linear functional on $X$ is bounded along bounded unconditional basic sequences simply because of absence of these sequences. So, take any unbounded functional $f:X\to\mathbb{R}$. Therefore, it makes sense to deal with the following:
@@
How to describe pairs of Banach spaces $X$ and $Y$ such that every operator $T:X\to Y$ bounded along normalized unconditional basic sequences of $X$ is bounded?
Although Question 1.2 is interesting and challenging, we are not going to study it in this paper.
@@
The example constructed by Storozhuk in [24] suggests that many Banach lattices (e.g., discrete Banach function spaces) do not belong this class.
The present paper deals with operators between OVSs and TVSs, so the order structure will be involved in the domain or/and codomain. We mainly focus on Banach spaces in the domain and, when suitable, Fréchet spaces. The case of Fréchet spaces in the domain occurs recently for compactly classified operators along disjoint bounded sequences in [19].
The present paper is organized as follows. Section 2 is devoted to order-to-topology continuous and order-to-topology bounded operators between OBSs (or OVSs) and TVSs. Several conditions providing topological boundedness of such operators are investigated (as well as the study undertaking of the Uniform Boundedness Principle for collectively qualified families of them). Section 3 is devoted to conditions which provide boundedness of Levi and Lebesgue operators. For example, Theorem 3.2 tells that each quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered TVS with a normal cone is topologically bounded, and Theorem 3.11 states that if an OBS in the domain has a closed generating normal cone then every $\sigma$-w-Lebesgue operator is bounded.
For further unexplained terminology and notation, we refer to [1, 3, 4].
@@ 2 On boundedness of some operators between ordered TVSs
Order-to-topology continuous and order-to-norm bounded operators have been studied recently (mostly in the Banach lattice setting) by many authors (see, for example, [5, 15, 16, 18, 19, 22, 23]). Some of their results were extended to operators from an OBS to a TVS in [14]. It is important to find further conditions on operators in ordered TVSs which provide automatic boundedness (see, e.g. [3, 4, 7, 13, 14]).
In this section, we continue the research initiated in [14, 15, 13, 16]. Propositions 2.3, 2.4, 2.5, and 2.6 supply with several conditions providing topological boundedness of order-to-topology continuous and order-to-norm bounded operators, and the Uniform Boundedness Principle for collectively qualified families of such operators.
Then, we adopt the definition of $\tau$-Levi and $\tau$-Lebesgue operators from [5, 20, 14] to the OVS-setting in the domain and investigate $\tau$-Lebesgue operators from an OBS to an NS. Lemma 3.10 states that each $\sigma$-w-Lebesgue operator is order-to-norm bounded whenever an OBS in the domain has normal cone. Theorem 3.11 states that if the domain has closed generating normal cone then every $\sigma$-w-Lebesgue operator is bounded. Theorem 3.3 gives conditions for ($\sigma$-)order to (weak) norm continuity of ($\sigma$-)(w-)Lebesgue operators.
This section is also aimed to introduce several notions and notation that will be used below.
Recall, that a net $(x_{\alpha})$ in an OVS $X$
@@
    relative uniform converges to $x$ (ru-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with $\pm(x_{\alpha}-x)\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$.
@@
    is eventually order bounded if there exist $[a,b]$ in $X$ and $\alpha_{0}$ such that $x_{\alpha}\in[a,b]$ for all $\alpha\geq\alpha_{0}$.
@@
    is o-Cauchy (resp., ru-Cauchy)) if the double net $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})$ o-converges (resp., ru-converges).
It is straightforward to see that o-convergent and ru-convergent (hence, o-Cauchy and ru-Cauchy) nets are eventually order bounded. In particular, o-convergent and ru-convergent (hence, o-Cauchy and ru-Cauchy) sequences are order bounded.
@@ 1. $a)$
    order bounded if $T[a,b]$ is order bounded for every order interval $[a,b]$ in $X$ ((the set of such operators is denoted by ${\cal L}_{ob}(X,Y))$.
@@ 2. $b)$
    order continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $Y$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ ((shortly $T\in{\cal L}_{oc}(X,Y))$.
@@ 3. $c)$
    ru-continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ in $Y$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ ((shortly $T\in{\cal L}_{rc}(X,Y))$.
Let $X$ be an OVS and $(Y,\tau)$ be a TVS. An operator $T\in{\cal L}(X,Y)$ is
@@ 1. $d)$
    order-to-topology continuous ((or $T\in{\cal L}_{o{\tau}c}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write ${\cal L}_{onc}(X,Y)$ when $Y$ is an NS)).
@@ 2. $e)$
    ru-to-topology continuous ((or $T\in{\cal L}_{r{\tau}c}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write ${\cal L}_{rnc}(X,Y)$ when $Y$ is an NS)).
@@ 3. $f)$
    order-to-topology bounded ((or $T\in{\cal L}_{o{\tau}b}(X,Y))$ if $T[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal L}_{onb}(X,Y)$ when $Y$ is an NS)).
Clearly, the sets ${\cal L}_{ob}(X,Y))$, ${\cal L}_{rc}(X,Y))$, ${\cal L}_{onc}(X,Y)$, etc., given by Definition 2.1 are vector spaces.
[…]
@@ 1. $a)$
    collectively order bounded ((or ${\cal T}\in\text{\bf L}_{ob}(X,Y)$)) whenever ${\cal T}[a,b]$ is order bounded for every $a,b\in X$.
Let $X$ be an OVS and $(Y,\tau)$ be a TVS. A subset $\cal T$} of ${\cal L}(X,Y)$ is
@@ 1. $b)$
    collectively order-to-topology continuous $(or{\cal T}\in\text{\bf L}_{o{\tau}c}(X,Y)$)) if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((when $Y$ is an NS, we write ${\cal T}\in\text{\bf L}_{onc}(X,Y))$.
@@ 2. $c)$
    collectively ru-to-topology continuous $(or{\cal T}\in\text{\bf L}_{r{\tau}c}(X,Y)$)) if $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((we write ${\cal T}\in\text{\bf L}_{rnc}(X,Y)$ for an NS $Y)$.
@@ 3. $d)$
    collectively order-to-topology bounded $(or{\cal T}\in\text{\bf L}_{o{\tau}b}(X,Y)$)) if ${\cal T}[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ for an NS $Y)$.
@@ 1. $e)$
    collectively continuous $(or{\cal T}\in\text{\bf L}_{c}(X,Y)$)) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
@@ 2. $f)$
    collectively bounded $(or{\cal T}\in\text{\bf L}_{b}(X,Y)$)) if ${\cal T}U$ is $\tau$-bounded whenever $U$ is $\xi$-bounded.
The $\sigma$-versions of Definition 2.1 b–e) and Definition 2.2 b), c), and e) are obtained via replacing nets by sequences (with the corresponding notations ${\cal L}_{oc}^{\sigma}(X,Y)$, ${\cal L}_{o{\tau}c}^{\sigma}(X,Y)$, $\text{\bf L}_{c}^{\sigma}(X,Y)$, etc.).
In many cases, collectively order-to-topology bounded sets agree with collectively ru-to-topology continuous ones.
@@
([14, Theorem 2.1]) We have $\text{\bf L}_{o{\tau}b}(X,Y)\subseteq\text{\bf L}_{r{\tau}c}(X,Y)$, whenever $X$ is an OVS and $Y$ is a TVS. If, additionally, $X_{+}-X_{+}=X$ then $\text{\bf L}_{o{\tau}b}(X,Y)=\text{\bf L}_{r{\tau}c}(X,Y)$.
For the single operator, it follows that every order-to-topology bounded operator from an OVS $X$ to a TVS $Y$ is ru-to-topology continuous (and if, additionally, $X_{+}$ is generating, ${\cal L}_{o{\tau}b}(X,Y)={\cal L}_{r{\tau}c}(X,Y)$).
The following theorem is a topological version of [16, Theorem 2.1].
@@
([14, Theorem 2.3]) We have $\text{\bf L}_{o{\tau}c}(X,Y)\subseteq\text{\bf L}_{o{\tau}b}(X,Y)$, whenever $X$ is an Archimedean OVS, $X_{+}-X_{+}=X$, and $Y$ a TVS.
In particular, every order-to-topology continuous operator from an Archimedean OVS with a generating cone to a TVS is order-to-topology bounded.
The following proposition is a partial extension of [15, Theorem 2.1] and [16, Theorem 2.8].
@@
([14, Theorem 2.5]) We have $\text{\bf L}_{o{\tau}b}(X,Y)\subseteq\text{\bf L}_{b}(X,Y)$, whenever $X$ is an OBS with a closed generating cone, and $Y$ is a TVS.
It follows that every collectively order-to-norm bounded set of operators from an OBS with a closed generating cone to an NS is uniformly bounded. In particular, each order bounded operator from an OBS with a closed generating cone to an ONS with a normal cone is continuous. Following [14], we collect few more consequences of Proposition 2.5 in the next proposition.
Proposition 2.6.
Let $X$ be an OBS with a closed generating cone.

### 16 Apr 2026 v3

# On automatic boundedness of operators from ordered to topological vector spaces

@@ Abstract
We study continuity and topological boundedness of order-to-topology bounded and order-to-topology continuous operators from ordered vector spaces to topological vector spaces. The Levi operators and Lebesgue operators are investigated.
@@ 11 Sobolev Institute of Mathematics, Novosibirsk, Russia
22 Southern Mathematical Institute of the Vladikavkaz Scientific Center of the Russian Academy of Sciences, Vladikavkaz, Russia
Keywords: ordered vector space, topological vector space, Banach space, Fréchet space, order-to-topology bounded operator
@@ 1 Introduction and preliminaries
Automatic boundedness of operators enjoying some additional properties plays significant role in analysis. For example, a nice theory exists on automatic continuity of algebra homomorphisms (see, e.g. [7, 8]). Another case is the automatic continuity of positive operators from Banach lattices to normed lattices [3, 4, 6]. Apparently, importance of boundedness of operators from a Banach space to a normed space lies in the Uniform Boundedness Principle for families of these operators. The present paper is devoted to the question how far one can go with other kinds of operators from a Banach space to a topological vector space. This question has attracted an attention in several recent papers (see, for example, [5, 11, 12, 13, 15, 14, 17, 18, 21, 22, 24]).
Throughout, vector spaces are real, and operators are linear. We abbreviate a topological vector space (normed space, locally convex space, ordered vector space, ordered normed space, ordered Banach space) by TVS (NS, LCS, OVS, ONS, OBS respectively). Symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net such that $\inf\limits_{\alpha}x_{\alpha}=0$; $a,b]=\{x\in X:a\leq x\leq b\}$ for an order interval; ${\cal L}(X,Y)$ for the space of operators between vector spaces $X$ and $Y$; $B_{X}(a,r)=\{x\in X:p(x,a)\leq r\}$ for the closed ball centered at $a\in X$ of radius $r\geq 0$ in a metric space $(X,p)$; and $B_{X}:=B_{X}(0,1)$, where $X$ is a NS. It is proved in [13, Theorem 2.1] that each order-to-norm bounded family of operators from a Banach lattice to an NS is equi-continuous. In [15, Theorem 2.1], this result is extended to the OBS setting.
Let $T:X\to Y$ be an operator from an NS $X$ to a TVS $(Y,\tau)$. In absence of additional structures in $X$, one cannot say much about conditions providing topological boundedness/continuity of $T$. For example, using of finite sets gives no proper subclass, because $T$ always carries finite sets into $\tau$-bounded sets. On the other hand, usage of compactness returns us to bounded operators. Indeed, if $T$ is $\tau$-bounded then $T$ takes compact sets into bounded sets. On the other hand, if $T$ is not $\tau$-bounded then there exist a circled $U\in\tau(0)$ and a sequence $(x_{n})$ in $B_{X}$ such that $Tx_{n}\notin n^{2}U$ for all $n\in\mathbb{N}$. In this case, the image $\big\{T(\frac{x_{n}}{n})\big\}_{n=1}^{\infty}\cup\{0\}$ of the compact set $\big\{\frac{x_{n}}{n}\big\}_{n=1}^{\infty}\cup\{0\}$ under $T$ is not $\tau$-bounded.
The case of a Hilbert space in the domain was investigated in papers [17, 18], where it is proved in [18, Theorem 2.1] (see also [17, Lemma 1.2]) that:
@@
Let us drop requirements of additional structures (e.g., lattice operations, order, or inner products) on a space in the domain and consider:
@@
The ‘‘conditional part’’ of Question 1.1 is open even when $X$ is reflexive, whereas the ‘‘unconditional part’’ has a negative answer due to the example of a reflexive Banach space with no unconditional basic sequence constructed by Gowers and Maurey [20] (every linear functional on this space is bounded along bounded unconditional basic sequences simply because of absence of such sequences, so, we take any unbounded functional on this space). Therefore, it makes sense to deal with the following questions.
@@
Under which conditions on Banach spaces $X$ and $Y$ every operator $T:X\to Y$ bounded along normalized unconditional basic sequences of $X$ is bounded?
@@
An example, recently constructed by Storozhuk in [23] suggests that many of Banach lattices (e.g., infinite dimensional discrete Banach function spaces) do not belong to this class.
The present paper deals with operators between OVSs and TVSs, so the order structure will be involved in the domain or/and codomain. We mainly focus on Banach spaces in the domain and (when suitable) on Fréchet spaces. The case of Fréchet spaces in the domain occurred recently in study of compactly classified operators along disjoint bounded sequences [18].
This paper is organized as follows. In Section 2, we study order-to-topology continuous/bounded operators between OBSs (or, also, OVSs) and TVSs. Conditions providing continuity of such operators are investigated. Section 3 is devoted to conditions providing boundedness/continuity of the Levi operators and the Lebesgue operators. For instance, Theorem 3.3 tells us that each quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered TVS with a normal cone is topologically bounded. Theorem 3.12 states that every $\sigma$-w-Lebesgue operator from an OBS with a closed generating normal cone to an NS is bounded.
For further unexplained terminology about TVSs and OVSs, we refer the reader to [1, 3, 4].
@@ 2 On boundedness of some operators between ordered TVSs
Order-to-topology continuous and order-to-norm bounded operators have been studied recently (mostly in the Banach lattice setting) by several authors (see, e.g., [5, 13, 15, 17, 18, 21, 22]). Some of their results were extended to operators from an OBS to a TVS in [12]. It is important to find further conditions on operators in ordered TVSs which provide automatic continuity/boundedness.
In this section, we continue the research initiated in [12, 13, 11, 15]. Propositions 2.3, 2.4, and 2.5, provides several conditions for topological boundedness of order-to-topology continuous and of order-to-norm bounded operators, and the Uniform Boundedness Principle for collectively qualified families of such operators.
Then, we study $\tau$-Levi and $\tau$-Lebesgue operators following [5, 19, 12] in more general OVS-setting in the domain. Lemma 3.11 states that each $\sigma$-w-Lebesgue operator is order-to-norm bounded whenever an OBS in the domain has normal cone. Theorem 3.12 states that, if the domain has closed generating normal cone, then every $\sigma$-w-Lebesgue operator is bounded. Theorem 3.4 gives conditions for ($\sigma$-)order to (weak) norm continuity of ($\sigma$-)(w-)Lebesgue operators.
We need several further notions and notation. Recall that a net $(x_{\alpha})$ in an OVS $X$
@@
    relative uniform converges to $x$ (ru-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{ru}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $\pm(x_{\alpha}-x)\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$.
@@
    is eventually order bounded if there exist $a,b\in X$ and $\alpha_{0}$ with $x_{\alpha}\in[a,b]$ for $\alpha\geq\alpha_{0}$.
@@
    is o-Cauchy (resp., ru-Cauchy) if $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})\operatorname{\xrightarrow[]{o}}0$ (resp., $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})\operatorname{\xrightarrow[]{ru}}0$).
It is straightforward to see that every o-convergent and every ru-convergent (hence, o-Cauchy and ru-Cauchy) net is eventually order bounded. In particular, o-convergent and ru-convergent (hence, also o-Cauchy and ru-Cauchy) sequences are order bounded.
@@ 1. $a)$
    order bounded if $T[a,b]$ is order bounded for every $a,b]\subseteq X$ ((shortly, $T\in{\cal L}_{ob}(X,Y))$.
@@ 2. $b)$
    order continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $Y$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ ((shortly, $T\in{\cal L}_{oc}(X,Y))$.
@@ 3. $c)$
    ru-continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ in $Y$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ ((shortly, $T\in{\cal L}_{rc}(X,Y))$.
An operator $T$ from an OVS $X$ to a TVS $(Y,\tau)$ is
@@ 1. $d)$
    order-to-topology continuous $(T\in{\cal L}_{otc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write $T\in{\cal L}_{onc}(X,Y)$, when $Y$ is an NS)).
@@ 2. $e)$
    ru-to-topology continuous $(T\in{\cal L}_{rtc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write $T\in{\cal L}_{rnc}(X,Y)$, when $Y$ is an NS)).
@@ 3. $f)$
    order-to-topology bounded $(T\in{\cal L}_{otb}(X,Y))$ if $T[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write $T\in{\cal L}_{onb}(X,Y)$, when $Y$ is an NS)).
An operator $T$ from a TVS $(X,\xi)$ to a TVS $(Y,\tau)$ is
  1. $g)$
    continuous $(T\in{\cal L}_{c}(X,Y))$ if $Tx_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
  2. $e)$
    topologically-bounded $(T\in{\cal L}_{tb}(X,Y))$ if $TU$ is $\tau$-bounded for every $\xi$-bounded $U$ ((we write ${\cal L}_{b}(X,Y)$, when $Y$ and $Y$ are NSs)).
Evidently, the sets ${\cal L}_{ob}(X,Y)$, ${\cal L}_{onc}(X,Y)$, etc. given by Definition 2.1 are vector spaces.
We recall several notions on collectively qualified sets of operators (cf. [9, 10, 12, 13, 17, 18, 15]). A family ${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ of nets indexed by a directed set $A$ in a TVS $(X,\tau)$ collective $\tau$-converges to 0 (shortly, ${\cal B}\operatorname{\xrightarrow[]{c\tau}}0$) if, for each $U\in\tau(0)$, there exists $\alpha_{U}$ such that $x^{b}_{\alpha}\in U$ for all $\alpha\geq\alpha_{U}$ and $b\in B$.
@@ 1. $a)$
    equi-order bounded $({\cal T}\in\text{\bf L}_{ob}(X,Y))$, whenever ${\cal T}[a,b]$ is order bounded for every $a,b\in X$.
A set $\cal T$} of operators from an OVS $X$ to a TVS $(Y,\tau)$ is
@@ 1. $b)$
    equi-order-to-topology continuous $({\cal T}\in\text{\bf L}_{otc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((when $Y$ is an NS, we write ${\cal T}\in\text{\bf L}_{onc}(X,Y))$.
@@ 2. $c)$
    equi-ru-to-topology continuous $({\cal T}\in\text{\bf L}_{rtc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((we write ${\cal T}\in\text{\bf L}_{rnc}(X,Y)$, when $Y$ is an NS)).
@@ 3. $d)$
    equi-order-to-topology bounded $({\cal T}\in\text{\bf L}_{otb}(X,Y))$ if ${\cal T}[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal T}\in\text{\bf L}_{onb}(X,Y)$, when $Y$ is an NS)).
@@ 1. $e)$
    equi-continuous $({\cal T}\in\text{\bf L}_{c}(X,Y))$ if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
@@ 2. $f)$
    equi-topologically-bounded $({\cal T}\in\text{\bf L}_{tb}(X,Y))$ if ${\cal T}U$ is $\tau$-bounded, whenever $U$ is $\xi$-bounded ((we write ${\cal T}\in\text{\bf L}_{b}(X,Y)$, when $X$ and $Y$ are NSs)).
The sequential versions of Definition 2.1 b)–e), g) and Definition 2.2 b), c), and e) are obtained via replacing nets by sequences (with notations ${\cal L}_{oc}^{\sigma}(X,Y)$, ${\cal L}_{otc}^{\sigma}(X,Y)$, $\text{\bf L}_{c}^{\sigma}(X,Y)$, etc.).
In many cases, equi-order-to-topology bounded sets agree with equi-ru-to-topology continuous ones.
@@
([12, Theorem 2.1]) $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{rtc}(X,Y)$, whenever $X$ is an OVS and $Y$ is a TVS. If additionally $X_{+}$ is generating then $\text{\bf L}_{otb}(X,Y)=\text{\bf L}_{rtc}(X,Y)$.
For the single operator, it follows that every order-to-topology bounded operator from an OVS $X$ to a TVS $Y$ is ru-to-topology continuous, and if additionally $X_{+}$ is generating, ${\cal L}_{otb}(X,Y)={\cal L}_{rtc}(X,Y)$. The following proposition is a topological counterpart of [15, Theorem 2.1].
@@
([12, Theorem 2.3]) $\text{\bf L}_{otc}(X,Y)\subseteq\text{\bf L}_{otb}(X,Y)$, whenever $X$ is an Archimedean OVS with a generating cone, and $Y$ is a TVS.
In particular, every order-to-topology continuous operator from an Archimedean OVS with a generating cone to a TVS is order-to-topology bounded. The following proposition is a partial extension of [13, Theorem 2.1] and [15, Theorem 2.8].
@@
([12, Theorem 2.5]) $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{b}(X,Y)$, whenever $X$ is an OBS with a closed generating cone, and $Y$ is a TVS.
It follows that every equi-order-to-norm bounded set of operators from an OBS with a closed generating cone to an NS is uniformly bounded. In particular, each order bounded operator from an OBS with a closed generating cone to an ONS with a normal cone is continuous. Following [12], we also collect few more statements below.
Let $X$ be an ordered Banach space with a closed generating cone.